OpenStudy (anonymous):

What is the solution to the equation? 4x-26-3x=16-20 Help? First accurate answer gets medal

OpenStudy (alexandervonhumboldt2):

x=-4+26 x=22

OpenStudy (anonymous):

Alex is wrong

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

first combine like terms

Nnesha (nnesha):

both side you have combine like terms do that first

OpenStudy (anonymous):

7x=22 correct?

OpenStudy (alexandervonhumboldt2):

4x-3x=x

OpenStudy (alexandervonhumboldt2):

16-20+26=22

OpenStudy (alexandervonhumboldt2):

---->x=22

OpenStudy (alexandervonhumboldt2):

@Nnesha

OpenStudy (anonymous):

@Nnesha Is he right?

Nnesha (nnesha):

there is 4x -3x = ??

Nnesha (nnesha):